# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v1

@@ Abstract
Machine Learning, ICML, OOD, adversarial examples
@@ 1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD (Madry et al., 2017), which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts (Taori et al., 2020) but also for adversarial examples (Sehwag et al., 2019a) without degrading the generalizability of the underlying pre-trained classifier (Yang et al., 2021).
Current standardized benchmarks such as OpenOOD (Zhang et al., 2023) and RoboDepth (Kong et al., 2024) merely focus on natural distribution shifts and corruptions (Hendrycks & Dietterich, 2019). Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection, model uncertainty, open set recognition, and OOD detection methods.
OpenOOD benchmark suite (Yang et al., 2021) evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) (Hendrycks & Gimpel, 2016) and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories).
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O (Wang et al., 2022), ImageNet-O (Hendrycks et al., 2021), Texture (Cimpoi et al., 2014), and iNaturalist (Huang & Li, 2021; Van Horn et al., 2018). OOD detection is a very quickly growing field due to the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data (Yang et al., 2022a; Cong & Prakash, 2022). Simple post-hoc methods like KNN (Sun et al., 2022) are highlighted maintaining good performance on toy-datasets (e.g. MNIST (Deng, 2012), CIFAR-10 or CIAR-100 (Krizhevsky et al., 2009), and also show outstanding performance on the more realistic dataset like ImageNet (Deng et al., 2009) according to (Yang et al., 2022a). These experiments do neglect the adversarial robustness of “state-of-the-art” detectors and the real-world capabilities are questionable as past studies had shown (Sehwag et al., 2019a; Song et al., 2020; Chen et al., 2020; Salehi et al., 2021). Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@ 2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate adversarial examples that will result in the misclassification of inputs at deep learning models (Biggio et al., 2013). It is possible to distinguish between two types of attack: black-box attacks (Zheng et al., 2023), where the classifier is queried, and white-box attacks (Carlini & Wagner, 2017b), where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. They try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ (Croce et al., 2020), to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by (Goodfellow et al., 2014) maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y)),$ where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by (Kurakin et al., 2018), using a reduced step size $\alpha$: ${\bm{x}}^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by (Madry et al., 2017), which is commonly considered a standard attack for the evaluation of model robustness (Liu et al., 2023a). Masked PGD (mPGD) (Xu et al., 2023b) is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
Lastly, DeepFool (DF) (Moosavi-Dezfooli et al., 2016) assumes that the network’s decision boundary is linear, even though in reality, it may not be. It aims to find the minimal perturbation, corresponding to the orthogonal projection onto the hyperplane.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy (Zhang et al., 2023). Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data (Zhang et al., 2023). In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE (Xu et al., 2023a). In contrast to the previous activation shaping method, ASH (Djurisic et al., 2022), that involves pruning and scaling of activations, SCALE demonstrates state-of-the-art results are solely derived from scaling.
OOD detection has a more expansive scope when compared to anomaly detection or open-set recognition (OSR) (Scheirer et al., 2012). Anomaly detection is concerned with the identification of rare deviations within a single distribution. OSR addresses the issue of unknown classes during inference. OOD detection methods aim to identify any test sample that deviates from the training data distribution (Zhang et al., 2023).
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR (Gillert & von Lukas, 2021) or uncertainty estimation (Schwaiger et al., 2020). Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation (Ledda et al., 2023) differ from the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE (Wilson et al., 2023).
@@ 2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors, but the task of providing comprehensive defense (Carlini et al., 2019) against unknown threats is challenging. Every defense mechanism can be circumvented at some point (Carlini & Wagner, 2017a). Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD adversarial examples that reveal a severe limitation of current open-world learning frameworks (Sehwag et al., 2019b; Azizmalayeri et al., 2022). Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD adversarial examples (Azizmalayeri et al., 2022). In the past years, various defensive techniques (Wu et al., 2023) and combinations of them have surfaced to combat threats such as adversarial training (Madry et al., 2017; Wang et al., 2023; Bai et al., 2024), gradient masking for obfuscation (Papernot et al., 2017), input transformations such as input purification (Nie et al., 2022; Lin et al., 2024). However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms (Tramer et al., 2020). According to (Croce et al., 2022), optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
OOD detectors have benefited from insights in the adversarial machine learning (AML) field, but still lack comprehensive defense against unknown threats. There are adversarial training based methods, e.g. ALOE (Chen et al., 2020), OSAD (Shao et al., 2020), or ATOM (Chen et al., 2021). A discriminator-based method, ADT (Azizmalayeri et al., 2022), significantly outperforms previous methods by addressing their vulnerabilities to strong adversarial attacks. More recently, the post-hoc method SAFE (Wilson et al., 2023) leverages the most sensitive layers in a pre-trained classifier through targeted input-level adversarial perturbations. To this end, “adversarial robust” OOD detection methods lag, being a comprehensive defense against unknown and adaptive threats remains an intricate challenge.
@@ 3 ROBUSTNESS DEFINITION
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to (Karunanayake et al., 2024). In this work, we focus on unified robustness, which belong to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ (Moreno-Torres et al., 2012). An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$. Table 4 in Appendix A gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, ATOM (see Section 2.3), aim to be adversarial robust. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD (Zhang et al., 2023) or RoboDepth (Kong et al., 2024) do not include unified robustness in their benchmarks at the moment. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution (Hendrycks et al., 2021), where OOD detection in large-scale semantic space has attracted increasing attention (Hendrycks et al., 2019), see Appendix A. Some OOD datasets have issues, where ID classes are part of the OD dataset (Bitterwolf et al., 2023). Recently, (Yang et al., 2023) found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that state-of-the-art OOD detectors are more sensitive to covariate shift and the advances in semantic shift detection are minimal. The investigation of adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection (Yang et al., 2021). The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let’s define an attention map ${\bm{A}}$ (Guo et al., 2022): The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM (Selvaraju et al., 2017) to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ (Rieger & Hansen, 2020).
@@ 4 EXPERIMENTS
We extend the OpenOOD framework (Zhang et al., 2023) to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors. As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox (Rauber et al., 2017); and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 (He et al., 2016a), ResNet-50 (He et al., 2016a), and Swin-T (Liu et al., 2021). The datasets are CIFAR-10 & CIFAR-100 (Krizhevsky et al., 2009), ImageNet-1K (Deng et al., 2009) and its variant ImageNet-200 with just 200 classes.

### v2

@@ Abstract
Keywords Adversarial examples  $\cdot$ OOD  $\cdot$ post-hoc detectors
@@ 1 INTRODUCTION
Adversarial robustness in the context of out-of-distribution (OOD) detection refers to the ability of a detector to correctly identify OOD samples even when they have been adversarial perturbed to evade a deep neural network (DNN). Evasion attacks, e.g. PGD [1], which are designed to fool deep learning classifiers, are difficult to spot as an outlier for OOD detectors. To prevent errors in real-world applications, it is crucial to detect OOD cases, not only for natural distribution shifts [2] but also for adversarial examples [3] without degrading the generalizability of the underlying pre-trained classifier [4].
Current standardized benchmarks such as OpenOOD [5] and RoboDepth [6] merely focus on natural distribution shifts and corruptions [7]. Especially OpenOOD aims to make a fair comparison across methods initially developed for anomaly detection, model uncertainty, open set recognition, and OOD detection methods.
OpenOOD benchmark suite [4] evaluates methods on semantic shift (e.g., samples that are semantically different from the training data, representing truly novel or unseen concepts.) [8] and covariate shift (e.g. samples that come from a different distribution than the training data, but still belong to the same semantic categories).
@@
Current OOD detection methods, as some listed in Table 1, achieve outstanding results on prominent OOD benchmarks, such as the OpenImage-O [9], ImageNet-O [10], Texture [11], and iNaturalist [12, 13]. OOD detection is a very quickly growing field due to the number of methods added to OpenOOD. More specific post-hoc methods with their plug-and-play capabilities on pre-trained classifiers are more flexible and scalable compared to methods that require full retraining on new OOD data [14, 15]. Simple post-hoc methods like KNN [16] are highlighted maintaining good performance on toy-datasets (e.g. MNIST [17], CIFAR-10 or CIAR-100 [18], and also show outstanding performance on the more realistic dataset like ImageNet [19] according to [14]. These experiments do neglect the adversarial robustness of “state-of-the-art” detectors and the real-world capabilities are questionable as past studies had shown [3, 20, 21, 22]. Adversarial examples remain challenging because they share the same semantics as the training data, but aim to modify the classifier’s output.
@@ 2.1 Evasion Attacks Crafting Inliers
The objective of evasion attacks is to generate adversarial examples that will result in the misclassification of inputs at deep learning models [23]. It is possible to distinguish between two types of attack: black-box attacks [24], where the classifier is queried, and white-box attacks [25], where the network is under the attacker’s complete control. A white-box threat model is strictly stronger. They try to find the smallest possible perturbation, often imperceptible to humans, to manipulate the model’s decision boundaries.
More formally, for an input ${\bm{x}}$ with the ground-truth label $y$, an adversarial example ${\bm{x}}^{\prime}$ is crafted by adding small noise $\bm{\delta}$ to ${\bm{x}}$ such that the predictor model loss ${\bm{J}}({\bm{x}}^{\prime},y)$ is maximized. The $L^{p}$ norm of the adversarial noise should be less than a specified value $\epsilon$, i.e., $\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\leq\epsilon$, e.g., $\epsilon=8/255$ [26], to ensure that the image does not change semantically. The attack method Fast Gradient Sign Method (FGSM) by [27] maximizes the loss function in a single step by taking a step towards the sign of the gradient of ${\bm{J}}({\bm{x}},y)$ w.r.t. to ${\bm{x}}$: ${\bm{x}}^{\prime}={\bm{x}}+\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime},y)),$ where the noise meets the $L^{\infty}$ norm bound $\epsilon$. Furthermore, this approach can be applied iteratively, as shown by [28], using a reduced step size $\alpha$: ${\bm{x}}^{\prime}_{0}={\bm{x}},\;{\bm{x}}^{\prime}_{t+1}={\bm{x}}^{\prime}_{t}+\alpha\epsilon\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}^{\prime}_{t},y))$, where in each step, the perturbation should be constrained within the $L^{\infty}$ ball of radius $\epsilon$. This constraint is a characteristic of the Projected Gradient Descent (PGD) attack proposed by [1], which is commonly considered a standard attack for the evaluation of model robustness [29]. Masked PGD (mPGD) [30] is a variant of the PGD attack that restricts perturbations to a specific region within an image. The PGD attack is extended towards a patch: ${\bm{x}}_{t+1}^{\prime}=\text{Clip}_{{\bm{x}}^{\prime}}\left({\bm{x}}_{t}^{\prime}+\alpha\cdot\text{sign}(\nabla_{{\bm{x}}}{\bm{J}}({\bm{x}}_{t}^{\prime},y,\bm{\theta})[patch]),\epsilon\right).$ In this context, the term “patch” refers to the region specified as $[x:x+h,y:y+w]$, where $[x,y,h,w]$ are the provided patch’s coordinates and dimensions.
Lastly, DeepFool (DF) [31] assumes that the network’s decision boundary is linear, even though in reality, it may not be. It aims to find the minimal perturbation, corresponding to the orthogonal projection onto the hyperplane.
@@ 2.2 Advantages of Post-Hoc OOD Detectors
Post-hoc OOD detection methods, which make use of specific layers of the pre-trained classifier, have been demonstrated to outperform retraining-based approaches, thereby underscoring their empirical efficacy [5]. Their plug-and-play nature allows seamless integration with pre-trained models, without necessitating alterations to the training procedure or access to original training data [5]. In Table 1 are 16 post-hoc detectors listed, which also shows if a method uses features, logits, or the probabilities of the pre-trained model. Post-hoc methods underline their simplicity to outperform others on natural distribution shifts datasets by being as lightweight as possible. The latest post-hoc OOD detector in Table 1 is SCALE [32]. In contrast to the previous activation shaping method, ASH [33], that involves pruning and scaling of activations, SCALE demonstrates state-of-the-art results are solely derived from scaling.
OOD detection has a more expansive scope when compared to anomaly detection or open-set recognition (OSR) [34]. Anomaly detection is concerned with the identification of rare deviations within a single distribution. OSR addresses the issue of unknown classes during inference. OOD detection methods aim to identify any test sample that deviates from the training data distribution [5].
Moreover, post-hoc OOD methods can be augmented with other techniques, such as those employed in OSR [35] or uncertainty estimation [36]. Combining different techniques means also the post-hoc methods become more complex and thus the attack surface might become larger, i.e. attacks against uncertainty estimation [37] differ from the evasion attacks. It is not necessarily the case that a post-hoc method must be combined with other techniques. This demonstrates SAFE [38].
@@ 2.3 OOD Adversarial Detection
Ensuring the protection of deployed DL models is the aim of OOD detectors, but the task of providing comprehensive defense [39] against unknown threats is challenging. Every defense mechanism can be circumvented at some point [40]. Many OOD detectors can be easily evaded by slightly perturbing benign OOD inputs, creating OOD adversarial examples that reveal a severe limitation of current open-world learning frameworks [41, 42]. Even adversarial training-based defense methods, effective against ID adversarial attacks, struggle against OOD adversarial examples [42]. In the past years, various defensive techniques [43] and combinations of them have surfaced to combat threats such as adversarial training [1, 44, 45], gradient masking for obfuscation [46], input transformations such as input purification [47, 48]. However, attackers consistently adapt their adversarial attacks to the specific defense mechanisms [49]. According to [50], optimization-based defenses could be a promising future, because they can adapt during test-time towards the input.
OOD detectors have benefited from insights in the adversarial machine learning (AML) field, but still lack comprehensive defense against unknown threats. There are adversarial training based methods, e.g. ALOE [21], OSAD [51], or ATOM [52]. A discriminator-based method, ADT [42], significantly outperforms previous methods by addressing their vulnerabilities to strong adversarial attacks. More recently, the post-hoc method SAFE [38] leverages the most sensitive layers in a pre-trained classifier through targeted input-level adversarial perturbations. To this end, “adversarial robust” OOD detection methods lag, being a comprehensive defense against unknown and adaptive threats remains an intricate challenge.
@@ 3 ROBUSTNESS DEFINITION
The robustness definition in the field of OOD detectors has been ambiguous when it comes to attack methods. There are two categories of adversarial examples. The first one merely attacks the underlying pre-trained classifier and the second one aims to fool the OOD detector itself. Adversarial robustness can be considered in a classifier (Unified Robustness) or the OOD detector (Robust OOD Detection) according to [53]. In this work, we focus on unified robustness, which belong to the covariate shift. For image classification, a dataset $\mathcal{D}=\{({\bm{x}}_{i},y_{i});{\bm{x}}_{i}\in{\mathcal{X}},y_{i}\in{\mathcal{Y}}\}$ sample from a training distribution $\hat{P}_{\rm{data}}({\bm{x}},y)$ is used to train some classifier $C:{\mathcal{X}}\rightarrow{\mathcal{Y}}$. In real-world deployments, distribution shift occurs when classifier $C$ receives data from test distribution $\hat{P}_{\rm{test}}({\bm{x}},y)$ where $\hat{P}_{\rm{data}}({\bm{x}},y)\neq\hat{P}_{\rm{test}}({\bm{x}},y)$ [54]. An OOD detector is a scoring function $s$ that maps an image ${\bm{x}}$ to a real number $\mathbb{R}$ such that some threshold $\tau$ arrives at the detection rule $f({\bm{x}}):\text{ID if }s({\bm{x}})\geq\tau\text{, OOD otherwise}$. Table 4 in Appendix A gives an overview of several detectors as well as considered ID and OOD datasets and model architectures. The ImageNet-1K dataset together with ResNet-50 architecture has become standard for ID. Popular OD datasets are iNaturalist, SUN, Places, and Textures. Some OOD detectors, such as ALOE, OSAD, ADT, ATOM (see Section 2.3), aim to be adversarial robust. They usually take evading attacks (see Section 2.1) such as FGSM or PGD as OOD. These computationally expensive methods only show empirical results on the small-scale CIFAR-10. A robust OOD detector is built to distinguish whether a perturbed input is OOD. Standardized OOD benchmark frameworks, i.e. OpenOOD [5] or RoboDepth [6] do not include unified robustness in their benchmarks at the moment. Consequently, both frameworks, give a false sense of encompassing open-world capabilities. They focus on natural distribution [10], where OOD detection in large-scale semantic space has attracted increasing attention [55], see Appendix A. Some OOD datasets have issues, where ID classes are part of the OD dataset [56]. Recently, [57] found a clean semantic shift dataset that minimizes the interference of covariate shift. The experiments show that state-of-the-art OOD detectors are more sensitive to covariate shift and the advances in semantic shift detection are minimal. The investigation of adversarial examples could gain insights into understanding the covariate shift and towards a generalized OOD detection [4]. The difference between benign ${\bm{x}}$ and the attacked counterpart ${\bm{x}}^{\prime}$ is the different attention of the pre-trained classifier $C$ per sample. Let’s define an attention map ${\bm{A}}$ [58]: The input image is passed through a classifier $C$ to obtain a feature map ${\bm{F}}$. A possible tool could be Grad-CAM [59] to investigate the attention change between benign ${\bm{x}}$ and adversarial example ${\bm{x}}^{\prime}$ on ${\bm{F}}$ [60].
@@ 4 EXPERIMENTS
We extend the OpenOOD framework [5] to consider adversarial attacks. We attack the pre-trained classifiers on the corresponding test sets and evaluate 16 post-hoc OOD detectors. As attack methods, we choose FGSM(-$L^{\infty}$), PGD(-$L^{\infty}$), DF(-$L^{2}$) from FoolBox [61]; and the mPGD-($L^{\infty}$). The attacked models are ResNet-18 [62], ResNet-50 [62], and Swin-T [63]. The datasets are CIFAR-10 & CIFAR-100 [18], ImageNet-1K [19] and its variant ImageNet-200 with just 200 classes.